OpenStudy (anonymous):

how to find the LCD of 12, 9, 60, and 45?

OpenStudy (mathhelp346):

multiply them all

OpenStudy (anonymous):

you meant LCM=least common multiple...then it would be 3

OpenStudy (anonymous):

no, LCD

OpenStudy (anonymous):

LCm and LCD are different......from what i saw when i googled it

OpenStudy (anonymous):

but you don't have any denominators in the given information

OpenStudy (ybarrap):

12 = 2^2×3 (3 prime factors, 2 distinct) 9 = 3^2 (2 prime factors, 1 distinct) 60 = 2^2×3×5 (4 prime factors, 3 distinct) 45 = 3^2×5 (3 prime factors, 2 distinct) LCM = get all unique prime factors, multiply their highest powers: LCM = 2^2 * 3^2 * 5 = 180 LCD is the LCM for denominators for rational numbers (e.g. fractions) Ans. 180